OpenStudy (anonymous):

When is the function concave up/concave down. Already found Inflection points?

OpenStudy (anonymous):

(A). f(x)=x^3/(x^2-1) Inflection point x=0 (B). f(x)=x^4/(x^2-1) No inflection point

OpenStudy (campbell_st):

ok... stationary points are found in the 1st derivative. the nature of the stationary point is x = a f''(a) > 0 then a minimum occurs f"(a) = 0 a horizontal pt of inflexion. f"(a) < 0 a maximum occurs

OpenStudy (anonymous):

ok..

OpenStudy (campbell_st):

you also need to test either side of the possible point of inflexion to find a change of sign.

OpenStudy (anonymous):

what i usually did was plug numbers on either side of the inflection point i.e -1 and 1 (from part a) into the second derivative. but then the answer is really confusing

OpenStudy (campbell_st):

for A its a horizontal pt of inflexion B is correct.

OpenStudy (campbell_st):

well you'll find x = 0 is a solution to f'(x) = 0... which means its a stationary point... but you don't know what type (max, min or pt of inflexion) so use the 2nd derivative test... and you'll find f"(0) = 0 so its a horizontal point of inflexion (special case) checking either side of the solution to f"(x) = 0 can be used.

OpenStudy (anonymous):

i know how to find the inflection point. I just want to know where the function is concave up or down. for the (b) part there are no inflection points. so i don't know how to proceed.

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

there is a local max and min... but you need to find those values from the 1st derivative I got \[f'(x) = \frac{x^2(x^2 -3)}{(x^2 - 1)^2}\] so stationary points occur at \[x = 0, \pm \sqrt3\] then \[f"(x) = \frac{2x(x^2 + 3)}{(x^2 - 1)^3}\] so x = 0 is a horizontal pt of inflexion since f"(0) = 0 \[f(\sqrt{3}) > 0 \] so a min \[f"(-\sqrt{3}) < 0 \] so a max